FIRST OBSERVATION OF THE SPIN-ORBIT INTERACTION BETWEEN THE X̃ ^{1}A_{1} AND THE ã ^{3}B_{1} STATES of SiH_{2} BY STIMULATED EMISSION PUMPING SPECTROSCOPY

\underline{H. ISHIKAWA}, Y. MURAMOTO, and N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

The energy separation and order between the triplet and the singlet electronic states have been one of the central issues of SiH_{2} radical from both chemical and spectroscopic points of view. However, any rotationally and/or vibrationally resolved observation of the triplet (ã ^{3}B_{1}) state has not yet been reported. Since the ã state is considered to be located ∼7000 cm^{−1} above the singlet (X̃ ^{1}A_{1}) state, it is expected that an effect of the singlet-triplet interaction appears among highly excited vibrational levels of the X̃ state. Thus, we have carried out the stimulated emission pumping (SEP) spectroscopy of SiH_{2} in the vibrational energy region up to 10000 cm^{−1}. In this paper, we will report an observation of a small but a definitive perturbation due to the singlet-triplet interaction in the SEP spectrum. We have observed fifty-one vibrational levels in the vibrational energy region of 4800–10000 cm^{−1}. Due to strong 1ν_{1}:2ν_{2} Fermi and 2ν_{1}:2ν_{3} Darling-Dennison resonances, vibrational levels having the same polyad quantum number, P = 2v_{1} + v_{2} + 2v_{3}, construct polyad structures. The vibrational levels observed belong to polyads of P = 5 − 10. In the case of P ≤ 9, all the vibrational energies observed were fitted very well by the effective Hamiltonian model in which the above resonances were considered. In the case of the P = 10 polyad, however, an unexpected splitting of the band was observed. It was confirmed that this splitting is due to the spin-orbit interaction between the X̃ and the ã states based on the rotational dependence of this perturbation. The internal energy of the triplet state observed was about 9645 cm^{−1} measured from the (000) level of the X̃ state. This level is tentatively assigned as (030), based on the theoretical calculation^{a}. Details of the analysis will be discussed at the presentation.

aW. GABRIEL et al., Chem. Phys. 174, 45 (1993).